OpenStudy (anonymous):

PLEASE HELP! A real estate company borrowed $276,000 to develop some condos. It amortized the loan to pay monthly for 4 years at 7% interest. Find the amount of each monthly payment.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

You'll use the formula \[\Large P = \frac{i*L}{1-(1+i)^{-n}}\] P = monthly payment i = rate per period L = amount loaned out n = number of periods

jimthompson5910 (jim_thompson5910):

P is unknown and that's what we want to find out i = 0.07/12 = 0.0058333 approximately L = 276,000 n = 4*12 = 48 months

jimthompson5910 (jim_thompson5910):

plug those values into the formula and evaluate

OpenStudy (anonymous):

ok im doing it now

OpenStudy (anonymous):

wouldnt it be -48

OpenStudy (anonymous):

i got the answer thank you so much

jimthompson5910 (jim_thompson5910):

no it's not -48

jimthompson5910 (jim_thompson5910):

oh did you mean -n is equal to -48 ?

OpenStudy (anonymous):

yes